# Theory and Computation of Substructure Characteristic Modes

### v1

@@ IV-B MoM & T-matrix hybrid
The matrix $\mathbf{Z}+\mathbf{U}^{\mathrm{T}}_{4}\mathbf{T}_{\mathrm{\mathrm{b}1}}\mathbf{U}_{4}$ is the impedance matrix of a metallic body in the presence of a dielectric object, with the second term being interpreted as a contribution of matrix $\mathbf{T}_{\mathrm{\mathrm{b}1}}$ to the background Green’s function [16]. Characteristic decomposition of this modified impedance matrix gives a spectrum similar to Fig. 1, with the entire conducting region in the presence of a dielectric sphere [28]. On the other hand, if substructure modes solely excitable from region $\varOmega_{\mathrm{c}}$ are desired, then this matrix is further modified and decomposed according to (11). The resulting spectrum of modal significances, where the entire region $\varOmega_{\mathrm{b}}$ is considered a background, is shown in Fig. 4. The equivalence of this formulation to (1) is demonstrated in the same figure.
IV-C Infinite ground plane
@@
Within the scattering formulation (1), the problem is solved by adding an image of region $\varOmega_{\mathrm{c}}$ and an image of the incident field. This results in a total electric field with vanishing tangential components at the ground plane $\varOmega_{\mathrm{b}}$. The symmetry of the incident field eliminates half of the spherical waves. Specifically, only spherical waves exhibiting electric dipole moment normal to the ground plane and tangential magnetic dipole moment remain. For example, placing the ground plane $\varOmega_{\mathrm{b}}$ in the $xy$ plane and denoting $l,m$ the degree and azimuthal numbers, respectively, only TE spherical waves with even $l+m$ and TM spherical waves with odd $l+m$ remain. The equivalence of the impedance-based formulation and scattering formulation is presented in Fig. 5.
The practicality of using substructure characteristic modes is demonstrated by comparing spectra from Figs. 24 and 5. In all cases, the black resonance peak appearing between $1.5\,\mathrm{GHz}$ and $2\,\mathrm{GHz}$ can be identified but would be lost if full characteristic decomposition was made, cf., Fig. 1. This resonance belongs to a mode responsible for the PIFA operation, and substructure decompositions show how it is affected by a particular background.
@@ V Utilization of Arbitrary Full-Wave Solver
V-A Planar Inverted-F Antennas
@@
A PIFA of the same dimensions as in Fig. 1 is studied once more, now with the addition of two example configurations of the dielectric substrate. Modal significance data for these examples, along with the previously studied PEC-only model, are shown in Fig. 6.
The reference case (PIFA #1) is made solely of PEC and represented by black traces. This case serves only as a verification of the scattering dyadic matrix procedure [7]. The resulting data are indistinguishable from the curves in Fig. 2. The two cases involving dielectrics are also shown in Fig. 6 and include two configurations of a lossless dielectric material with relative permittivity $\varepsilon_{\mathrm{r}}=2.33$. One case (PIFA #2) is built with two thin (1.575 mm) dielectric layers backing the parallel conductive layers in the PIFA system. The other case (PIFA #3) includes a dielectric block spanning the entire space between the ground plane and the upper motif.
It can be seen that the eigen-traces of PIFA #1 and PIFA #2 are comparable, which is given by the fact that thin substrate minimizes the effect of the dielectrics, i.e., the effective permittivity between the ground plane and the upper motif is close to unity, leading to only a slight downward shift in the resonant frequency of the dominant characteristic mode. This is not the case with PIFA #3, which behaves both qualitatively and quantitatively differently. The modal significance maxima are considerably shifted towards lower frequencies due to the reduced wavelength within the dielectric. The shape of the trace associated with the dominant mode is also visually quite different.
@@
Figure 6: The first three substructure characteristic modes for different planar F antennas. The first model is taken from Fig. 2 and serves as a verification. The second model is built on a substrate with a thickness of 1.575 mm and relative permittivity $\varepsilon_{\mathrm{r}}=2.33$. The third model has a dielectric block of $\varepsilon_{\mathrm{r}}=2.33$ filling the space between the ground plane and the antenna. All models are lossless and treated with an iterative algorithm that employs surface equivalence method-of-moment from Altair FEKO [30]. The traces were adaptively refined to obtain smooth curves with relatively few samples.
V-B Simplified CubeSat Model
@@
[…]
The spectrum shown in Fig. 7 is dominated by only one mode, an observation agreeing well with the similar arrangement in Fig. 2. Considering the 20% relative size reduction of PIFA in Fig. 7 as compared with Fig. 2, the dominant mode resonates at comparable frequency. The background structure contains a large dielectric block with many internal resonances (approximately 80 cavity resonances between 0.1 GHz and 2.6 GHz). Nevertheless, all these resonances are filtered out by the substructure formulation (1). The only exception is the mode having an abrupt increase of modal significance around 1.6 GHz. This spike is still well-modeled by the iterative algorithm by finely sampling the frequency axis.
[…]
@@ VI Discussion
With the exception of the CubeSat example shown in Fig. 7, all examples are based on the same PIFA-like controllable region in the presence of varying backgrounds. Substantial differences between the resulting modal characteristics in each example clearly illuminate the high potential impact of background objects on modal performance. This is particularly noticeable in Fig. 6, where the inclusion of thin dielectric support layers leads to non-negligible changes in modal significance. If outputs from characteristic mode analyses are used in the design of antennas, it is therefore critical to fully model any background objects, including dielectrics, rather than using simplified background models and small perturbation approximations. The approach presented in this manuscript facilitates this rigorous analysis.
@@
The second generalization is the addition of ports, a situation depicted in panel (c). A possible scenario might be a substructure problem of Section II, where a port is added to controllable degrees of freedom [^2] If the port is considered part of the background, the system will be lossy, which is not considered in this paper.. Although proposals exist for evaluating characteristic modes on antennas loaded by ports [33], we stress here that the ports present no qualitative change for the scattering formulation. Furthermore, the scattering formulation suggests that substructure characteristic modes can be evaluated even for circuits [21]. For example, the scattering matrices $\mathbf{S}_{\mathrm{b}}$ might represent a circuit without a particular region, which will be considered the controllable part.
@@
Lastly, panel (d) of Fig. 8 combines all preceding scenarios. A possible application of this scenario is the analysis and optimization of minimum scattering antenna [34] by inspecting characteristic modes and their significance [35], or techniques involving generalized scattering matrix [36, 37]. In such a case, the substructure characteristic modes are not reachable via the Schur complement method [9], and their formulation will be challenging if background Green’s function is employed. Yet, the scattering formulation stays always the same, solely demanding to evaluate scattering matrices $\mathbf{S}$ and $\mathbf{S}_{\mathrm{b}}$ for a situation with and without the controllable part of the structure.
@@ VII Conclusion
The examples throughout the paper demonstrate the flexibility of the method through its application to a family of problems based on a PIFA-like antenna above a finite ground plane. The resulting data demonstrate equivalence between the proposed method and impedance-based formulations, and highlight the generality of the approach to variations involving large dielectric background media, infinite ground planes, and finite dielectric regions. Further discussion regarding its application to problems involving wave ports and acceleration using iterative algorithms points to several areas for continued research, with the general direction aiming toward the fast and efficient characteristic mode analysis of antennas and scatterers in highly complex, arbitrary material environments.
@@ Appendix A Matrix-Free Evaluation of Substructure Characteristic Modes
Characteristic modes can be evaluated efficiently using an iterative matrix-free algorithm [8]. Instead of having full, explicit knowledge of a matrix being decomposed, these algorithms rely on knowing the result of a matrix-vector multiplication of that matrix with an arbitrary vector. In the case of characteristic modes, the matrix itself is a scattering operator, while the matrix-vector multiplication represents the solution to a scattering problem for a particular excitation. Nevertheless, algorithms such as Arnoldi iteration used in [8] are not suitable for generalized eigenvalue problems, such as (1). Therefore, to evaluate substructure characteristic modes in a matrix-free fashion, the formulations (4), (6) and (7) are used. Furthermore, it is important to realize that only results of multiplications $\mathbf{T}\mathbf{x},\mathbf{T}_{\mathrm{b}}\mathbf{x},\mathbf{S}\mathbf{x},\mathbf{S}_{\mathrm{b}}\mathbf{x}$ are accessible via general purpose electromagnetic solvers. When adapting the algorithms described in [8], the following changes in the desired matrix-vector products are made
@@
An example of the procedure used to estimate substructure characteristic modes using transition matrices in matrix-free manner is sketched in Algorithm 1, where $\mathbf{M}=\mathbf{T}$, $\mathbf{M}_{\mathrm{b}}=\mathbf{T}_{\mathrm{b}}$, $\mathbf{A}=2\mathbf{T}_{\mathrm{b}}^{\mathrm{H}}\mathbf{T}+\mathbf{T}_{\mathrm{b}}^{\mathrm{H}}+\mathbf{T}$ abbreviates the matrix to be decomposed, and $\mathbf{P}=\mathbf{I}$ is an identity matrix. The algorithm is stopped when magnitude $\left|\mathbf{a}_{m+1}\right|$ is sufficiently small or relative changes in estimated eigenvalues $t_{n}$ are sufficiently small. Steps no. 5 and 6 are the solutions to scattering problems involving full structure and background, respectively, and can be obtained from any full-wave electromagnetic solver. The modified Gram–Schmidt procedure is used in Algorithm 1 to assure its stability [38].
Another possibility to evaluate substructure characteristic modes is to employ scattering dyadic matrices, accessible with arbitrary electromagnetic solver [7]. In this case, the matrices $\tilde{\mathbf{S}}$ and $\tilde{\mathbf{S}}_{\mathrm{b}}$ are defined as in [7, eq. (21)] and denoted here as scattering and background scattering dyadic matrices, respectively. The matrices are not transposed symmetric [7, eq. (4)], so the Algorithm 1 must be modified by settings $\mathbf{M}=\tilde{\mathbf{S}}$$\mathbf{M}_{\mathrm{b}}=\tilde{\mathbf{S}}_{\mathrm{b}}$, and with matrix $\mathbf{P}$ being indexing matrix containing zeros except positions $P_{pq}$ and $P_{pq}$ where pairs of the quadrature points are mapped as
$P_{pq}=P_{pq}=1\Longleftrightarrow\hat{\boldsymbol{r}}_{p}=-\hat{\boldsymbol{r}}_{q}.$ (18)

### v3

@@ IV-C MoM and T-matrix Hybrid
The matrix $\mathbf{Z}+\mathbf{U}^{\mathrm{T}}_{4}\mathbf{T}_{\mathrm{\mathrm{b}1}}\mathbf{U}_{4}$ is the impedance matrix of a metallic body in the presence of a dielectric object, with the second term being interpreted as a contribution of matrix $\mathbf{T}_{\mathrm{\mathrm{b}1}}$ to the background Green’s function [16]. Characteristic decomposition of this modified impedance matrix would lead to characteristic modes of the entire metallic body in the presence of a dielectric sphere [44]. On the other hand, if substructure modes solely excitable from region $\varOmega_{\mathrm{c}}$ are desired, then this matrix is further modified and decomposed according to (11). The resulting spectrum of modal significances, where the entire region $\varOmega_{\mathrm{b}}$ (the sphere together with the bottom metallic plate) is considered a background, is shown in Fig. 5. The equivalence of the modes of this formulation to (1) is demonstrated in the same figure. In the scattering formulation, the hybrid method [43] was instead used to obtain scattered field from the excitations $\mathbf{a}$ and, therefore, to obtain the corresponding transition matrices $\mathbf{T}$ and $\mathbf{T}_{\mathrm{b}}$.
IV-D Infinite Ground Plane
@@
Within the scattering formulation (1), the problem is solved by adding an image of region $\varOmega_{\mathrm{c}}$ and an image of the incident field. This results in a total electric field with vanishing tangential components at the ground plane $\varOmega_{\mathrm{b}}$. The symmetry of the incident field eliminates half of the spherical waves. Specifically, only spherical waves exhibiting electric dipole moment normal to the ground plane and tangential magnetic dipole moment remain. For example, placing the ground plane $\varOmega_{\mathrm{b}}$ in the $xy$ plane and denoting $l,m$ the degree and azimuthal numbers, respectively, only TE spherical waves with even $l+m$ and TM spherical waves with odd $l+m$ remain. The equivalence between modes produced by the impedance and scattering formulations is presented in Fig. 6.
The practicality of using substructure characteristic modes is demonstrated by comparing spectra from Figs. 25 and 6. In all cases, the black resonance peak appearing between $1.5\,\mathrm{GHz}$ and $2\,\mathrm{GHz}$ can be identified but would be lost if full characteristic decomposition was made, cf., Fig. 1. This resonance belongs to a mode responsible for the PIFA operation, and substructure decompositions show how it is affected by a particular background.
@@ V Utilization of Arbitrary Full-Wave Solver
V-A Planar Inverted-F Antennas (PIFA)
@@
A PIFA of the same dimensions as in Fig. 1 is studied once more, now with the addition of two example configurations of dielectric substrates. Modal significance data for these examples, along with the previously studied PEC-only model, are shown in Fig. 7. Five dominant modes are shown to emphasize the effect of the different dielectric configurations on detuning the fundamental resonance behaviors of the structure.
@@
Figure 7: The first five substructure characteristic modes for different planar inverted-F antennas. The first model is taken from Fig. 2 and serves as a verification. The second model is built on a substrate with a thickness of 1.575 mm and relative permittivity $\varepsilon_{\mathrm{r}}=2.33$. The third model has a dielectric block of $\varepsilon_{\mathrm{r}}=2.33$, filling the space between the ground plane and the antenna. All models are lossless. The controllable and background regions $\varOmega_{\mathrm{c}}$ and $\varOmega_{\mathrm{b}}$ are depicted by the yellow and blue colors in the top pane, respectively. Full lines belong to the iterative algorithm that employs surface equivalence method-of-moments from Altair FEKO [46]. The traces were adaptively refined to obtain smooth curves with relatively few frequency samples. The markers belong to solution via (1), where scattering matrices were obtained by a solver [48] combining surface and volumetric electric field integral equation.
For comparison, markers are used in the same figure to present results of (1), where the scattering matrices were obtained by a solver combining surface and volumetric electric field integral equation [48].
The reference case (PIFA #1) is made solely of PEC and represented by black traces. This case serves only as a verification of the scattering dyadic matrix procedure [7]. The resulting data are indistinguishable from the curves in Fig. 2. The two cases involving dielectrics are also shown in Fig. 7 and include two configurations of a lossless dielectric material with relative permittivity $\varepsilon_{\mathrm{r}}=2.33$. One case (PIFA #2) is built using thin dielectric layers, while the other case (PIFA #3) includes a dielectric block spanning the entire space between the ground plane and the upper motif.
It can be seen that the eigen-traces of PIFA #1 and PIFA #2 are comparable, which is given by the fact that thin substrate reduces the effect of the dielectrics, i.e., the effective permittivity between the ground plane and the upper motif is close to unity, leading to only a slight downward shift in the resonant frequency of the dominant characteristic mode. The case with PIFA #3 behaves similarly, but the modal significance maxima are shifted considerably towards lower frequencies due to the reduced wavelength within the dielectric. The frequency dependence of the trace associated with the dominant mode also differs significantly past the resonance.
By studying the detuning effects of different dielectric configurations, this example demonstrates just one potential analysis method afforded by the scattering formulation without restriction to a particular integral representation.
Figure 8: Characteristic current density and characteristic directivity pattern of the dominant characteristic mode of PIFA #3. The left panel corresponds to the maximum at frequency $f\approx 0.95\,\mathrm{GHz}$ and shows a maximum directivity $D\approx 2.1$. The right panel corresponds to the maximum at frequency $f\approx 1.3\,\mathrm{GHz}$ and shows a maximum directivity $D\approx 3.1$.
As presented in (15), once the eigenvector $\mathbf{a}_{n}$ is obtained from (1), the evaluation of other characteristic quantities is straightforward. A sample calculation of characteristic current density and characteristic directivity patterns is presented in Fig. 8 for the dominant mode of PIFA #3. The low-frequency figure shows a formation of an electric dipole-like radiation backed by the ground plane, still having considerable back radiation. At resonance of the dominant mode, the radiation diagram resembles the one used on fed PIFAs. The expansion of the current density into substructure characteristic modes on fed single resonance PIFA would be dominated by this mode.
V-B PIFA Mounted on an Electrically Large Platform
@@
The substructure characteristic modes of a PIFA mounted on an electrically large platform are evaluated in Fig. 9. The frame of dimensions $10\times 10\times 15$ cm3 (approximately 1.5U CubeSat format [49]) is formed by a PEC strip of width 1 cm, and is galvanically connected with the solid PEC top cover and left open on the underside. The interior is filled with a dielectric material with relative permittivity $\varepsilon_{\mathrm{r}}=3$. The PIFA of dimensions $w=5$ cm, $d=2.5\,$cm, and $h=1.25\,$cm is mounted on the top cover of the frame and, similarly as in Fig. 2, only this antenna region is considered controllable for modal analysis, see the inset in Fig. 9. The model is lossless and treated with an iterative algorithm that employs surface equivalence method-of-moment from Altair FEKO [46]. The traces were adaptively refined to obtain smooth curves with relatively few samples. Numerical solver AToM [48] was employed to independently verify the traces in a similar manner as in Fig. 7.
The spectrum shown in Fig. 9 is dominated by only one mode, which agrees well with the similar arrangement in Fig. 2. Considering the 20% relative size reduction of PIFA in Fig. 9 as compared with Fig. 2, the dominant mode resonates at comparable frequency. The background structure contains a large dielectric block with many internal resonances (approximately 80 cavity resonances between 0.1 GHz and 2.6 GHz). Most of these resonances are filtered out by the substructure formulation (1). The only exception is the mode having an abrupt increase of modal significance around 1.6 GHz. The reason is that this peak, similar to the wavy behavior of the red, green, and blue traces around 2.5 GHz, is a numerical artifact caused by a sensitive and imperfect cancellation between eigenvalues of $\mathbf{S}$ and $\mathbf{S}_{\mathrm{b}}$ matrices, cf., (1). This happens when the background region dominates the EM behavior of the entire structure, i.e., when the spectra of $\mathbf{S}$ and $\mathbf{S}_{\mathrm{b}}$ are similar. To be more specific and taking the explanation from Sec. IV-A, the sharp blue peak results from an imperfect cancellation of two eigenvalues $t_{n}$ (in the complex plane) which are almost identical for the full case and the background. In contrast, the dominant substructure mode represented by the black line in Fig. 9 results from a mode that only exists on the full structure and has no pair in the spectrum of the background. This numerical issue was not observed in the previous examples where the electrical size of the background was not sufficient to overshadow the scattering behavior of the controllable region.
Figure 9: The first five substructure characteristic modes of a PIFA mounted on a frame. The frame has dimensions $10\times 10\times 15$ cm3 and is made of PEC and has a width of 1 cm. The top cover is also made of PEC and forms a ground plane for a PIF. The PIFA has dimensions of $w=5$ cm, $d=2.5$ cm, $h=1.25$ cm, and is the only controllable region in the model. The frame is filled with dielectrics of relative permittivity $\varepsilon_{\mathrm{r}}=3$. The meaning of the solid and dashed lines is the same as in Fig. 7.
@@ VI Discussion
As an example of treating a complex material distribution,  Fig. 10 presents characteristic data for a setup similar to Fig. 5 but with a bi-isotropic medium described by constitutive relations
[…]
where $\varepsilon_{0}$, $\mu_{0}$, and $c_{0}$ are the vacuum permittivity, permeability, and speed of light, respectively. The evaluation of the characteristic spectrum solely from integral equation solvers presents a considerable difficulty for the use of the bi-isotropic Green’s function or the combined magneto-electric volume formulation. In contrast, the scattering-based formulation relying on (1) only demands the use of any electromagnetic solver that can solve two scattering problems. Here the problem is solved using the scattering formulation (1) and the hybrid MoM & T-matrix method described in Sec. IV-C as a benchmark. Note that the latter does not represent a fully MoM-based implementation of this problem, as the authors are unaware of any such implementations are available for this class of problem.
Figure 10: Modal significances $|t_{n}|$ computed for the same setup as in Fig. 5 but with a sphere made of bi-isotropic medium with relative permittivity $\varepsilon_{\mathrm{r}}=4$, relative permeability $\mu_{\mathrm{r}}=1$, electro-magnetic susceptibility $\chi^{\mathrm{em}}=1+\mathrm{j}$ and magneto-electric susceptibility $\chi^{\mathrm{me}}=\left(\chi^{\mathrm{em}}\right)^{\mathrm{H}}$. The presence of the real part of susceptibility $\chi^{\mathrm{em}}$ makes the system non-reciprocal. The impedance-based method relies on (11) and (17), while the scattering-based formulation relies on (1).
VI-A Impact on Antenna Design
With the exception of the PIFA mounted on conductive frame example shown in Fig. 9, all examples are based on the same PIFA-like controllable region in the presence of varying backgrounds. Substantial differences between the resulting modal characteristics in each example clearly illuminate the high potential impact of background objects on modal performance. This is particularly noticeable in Fig. 7, where the inclusion of thin dielectric support layers leads to non-negligible changes in modal significance. If outputs from characteristic mode analyses are used in the design of antennas, it is therefore critical to fully model any background objects, including dielectrics, rather than using simplified background models and small perturbation approximations. The approach presented in this manuscript facilitates this rigorous analysis.
VI-B Generalization to Other Substructure Problems
@@
The second generalization is the addition of ports, a situation depicted in panel (c). A possible scenario might be a substructure problem of Section II, where a port is added to controllable degrees of freedom [^4] If the port is considered part of the background, the system will be lossy, which is not considered in this paper.. Although proposals exist for evaluating characteristic modes on antennas loaded by ports [50], we stress here that the ports present no qualitative change for the scattering formulation. Furthermore, the scattering formulation suggests that substructure characteristic modes can be evaluated even for circuits [30]. For example, the scattering matrix $\mathbf{S}_{\mathrm{b}}$ might represent a circuit without a particular region, which is considered the controllable part.
@@
Lastly, panel (d) of Fig. 11 combines all preceding scenarios, ports still belonging to the controllable degrees of freedom. A possible application of this scenario is the analysis and optimization of minimum scattering antenna [51] by inspecting characteristic modes and their significance [52], or techniques involving generalized scattering matrix [53, 54]. In such a case, the substructure characteristic modes are not reachable via the Schur complement method [9], and their formulation will be challenging if background Green’s function is employed. Yet, the scattering formulation stays always the same, solely demanding to evaluate scattering matrices $\mathbf{S}$ and $\mathbf{S}_{\mathrm{b}}$ for a situation with and without the controllable part of the structure.
To provide a particular example of scenarios from Fig. 11c and Fig. 11d, consider two co-planar strips lying above a finite ground plane, see the inset of Fig. 12. The strips are in their centers connected to delta-gap ports with impedance $Z_{0}=73\,\mathrm{\Omega}$. The ground plane is considered as a background. The modal significances of substructure characteristic modes are shown in Fig. 12. The figure shows the case with ports being part of controllable degrees of freedom (corresponding to Fig. 11c) as well as the case when ports are missing (corresponding to Fig. 11b). It can be seen that the presence of ports has a considerable effect on the modal spectrum, particularly for modes with high current density at the location of the ports (green and black traces), which are effectively broadened by the inclusion of the ports. Two modes with zero current density at the position of the ports (red and blue) are unchanged between the models with and without ports. It is also important to notice that far fields attached to the case of Fig. 11c or Fig. 11d are orthogonal only when waves exiting the ports are also taken as a part of the far field. The sole radiation diagrams are not orthogonal as they are in the cases of Fig. 11a and Fig. 11b.
Figure 12: Modal significances $\left|t_{n}\right|$ computed for the two parallel strips (controllable part) in the presence of a ground plane (background). The strips are potentially connected to lumped ports (controllable part). The strips have length $l\approx 128\,\mathrm{mm}$ and width $l/20$. The distance between the dipole centers is $l/3$. The ground plane is a square of edge $23l/$ 2 and is positioned at the distance $l/5$ below the strips. All materials are perfect electric conductors.
VI-C Numerical Accuracy and Computational cost
The scattering formulation offers the ability to use arbitrary numerical tools when evaluating substructure characteristic modes. In this way, it allows for the judicious selection of numerical tools that can accurately and efficiently evaluate scattering from the particular structure and background being studied. As discussed in Section IV-B, certain problems benefit from the use of particular numerical methods, including situations where integral equations (and by proxy the existing Schur-complement approach to substructure modes) are the optimal choice for efficiency and accuracy.
Here, we reiterate that the scattering formulation opens the door for the analysis of substructure characteristic modes using other numerical methods and that the primary benefit of this extension is the ability to solve problems that are difficult or impossible to efficiently or accurately study using integral equations. Regardless of which method or numerical tool is selected, it is important to keep in mind that, ultimately, the accuracy of modal data relies on the accuracy of the underlying numerical modeling, be it scattering simulations or the construction of integral operators.
@@ VII Conclusion
The examples throughout the paper demonstrate the flexibility of the method through its application to a family of problems based on a PIFA-like antenna above a finite ground plane. The resulting data demonstrate equivalence between the modes from the proposed method with those from impedance-based formulations for problems where both are applicable, and highlight the generality of the approach to variations involving large dielectric background media, infinite ground planes, and finite dielectric regions. Further discussion regarding its application to problems involving wave ports and acceleration using iterative algorithms points to several areas for continued research, with the general direction aiming toward the fast and efficient characteristic mode analysis of antennas and scatterers in highly complex, arbitrary material environments.
@@ Appendix A Matrix-Free Evaluation of Substructure Characteristic Modes
Characteristic modes can be evaluated efficiently using an iterative matrix-free algorithm [8]. Instead of having full, explicit knowledge of a matrix being decomposed, these algorithms rely on knowing the result of a matrix-vector multiplication of that matrix with an arbitrary vector. In the case of characteristic modes, the matrix itself is a scattering operator, while the matrix-vector multiplication represents the solution to a scattering problem for a particular excitation. Nevertheless, algorithms such as Arnoldi iteration [55] used in [8] are not suitable for generalized eigenvalue problems, such as (1). Therefore, to evaluate substructure characteristic modes in a matrix-free fashion, the formulations (4), (6) and (7) are used. Furthermore, it is important to realize that only results of multiplications $\mathbf{T}\mathbf{x},\mathbf{T}_{\mathrm{b}}\mathbf{x},\mathbf{S}\mathbf{x},\mathbf{S}_{\mathrm{b}}\mathbf{x}$ are accessible via general purpose electromagnetic solvers. When adapting the algorithms described in [8], the following changes in the desired matrix-vector products are made
@@
An example of the procedure used to estimate substructure characteristic modes using transition matrices in matrix-free manner is sketched in Algorithm 1, where $\mathbf{M}=\mathbf{T}$, $\mathbf{M}_{\mathrm{b}}=\mathbf{T}_{\mathrm{b}}$, $\mathbf{A}=2\mathbf{T}_{\mathrm{b}}^{\mathrm{H}}\mathbf{T}+\mathbf{T}_{\mathrm{b}}^{\mathrm{H}}+\mathbf{T}$ abbreviates the matrix to be decomposed, and $\mathbf{P}=\mathbf{1}$ is an identity matrix. The algorithm is stopped when the magnitude $\left|\mathbf{a}_{m+1}\right|$ is sufficiently small or relative changes in estimated eigenvalues $t_{n}$ are sufficiently small. Steps no. 5 and 6 are the solutions to scattering problems involving full structure and background, respectively, and can be obtained from any full-wave electromagnetic solver. The modified Gram–Schmidt procedure is used in Algorithm 1 to assure its stability [55]. Algorithm 1 is, for simplicity, presented using matrices $\mathbf{A}_{m}$ and $\mathbf{P}_{m}$. However, it is important to note that storing these as full matrices is not required for evaluation of the eigenvalues [55, Ch. 10].
Another possibility to evaluate substructure characteristic modes is to employ scattering dyadic matrices, accessible with arbitrary electromagnetic solver [7]. In this case, the matrices $\tilde{\mathbf{S}}$ and $\tilde{\mathbf{S}}_{\mathrm{b}}$ are defined as in [7, eq. (21)] and denoted here as scattering and background scattering dyadic matrices, respectively. The matrices are not transposed symmetric [7, eq. (4)], so the Algorithm 1 must be modified by settings $\mathbf{M}=\tilde{\mathbf{S}}$$\mathbf{M}_{\mathrm{b}}=\tilde{\mathbf{S}}_{\mathrm{b}}$, and with matrix $\mathbf{P}$ being indexing matrix containing zeros except positions $P_{pq}$ and $P_{qp}$ where pairs of the quadrature points are mapped as
$P_{pq}=P_{qp}=1\Longleftrightarrow\hat{\boldsymbol{r}}_{p}=-\hat{\boldsymbol{r}}_{q}.$ (20)